# Elastic Data Transfer Optimization with Hybrid Reinforcement Learning

### v1

@@ II Related Work & Motivation
Trade-off Between Performance and Overhead: Both heuristic and supervised machine learning models estimate parameter values before data transfers begin, using fixed configurations throughout the entire process. This static approach prevents adaptation to changing network conditions during long-running transfers. Additionally, these solutions lack a back-off mechanism, meaning that large parameter values may overwhelm systems in various production environments. For example, if $x$ parallel TCP connections are required to fully utilize the available bandwidth, assigning $x$ connections to every new transfer becomes inefficient when multiple concurrent transfers are initiated. In a scenario with $N$ concurrent transfers, $Nx$ connections would be created, leading to diminished throughput per connection (by a factor of $N$), but significantly increased overhead (Figure LABEL:fig:cc_impacts). A more efficient alternative would be to allocate fewer connections, say, 10% of $x$ – to each new transfer, under the assumption that overlapping transfers or other processes will compete for I/O resources. For instance, Globus Online, a widely adopted data transfer service, allows tuning of three key parameters: pipelining, parallelism, and concurrency. However, in production environments, system administrators set these values heuristically and often very conservatively to minimize I/O and network contention. While this approach helps balance overhead with resource utilization, it frequently leaves networks underutilized. As a result, static configurations present a dilemma: either prioritize utilization at the cost of high overhead, or reduce overhead and risk underutilizing the network.
@@
Simplified Architecture Overwhelm Systems: The most recent studies also have implemented adaptive approaches using online black-box optimization algorithms such as Bayesian methods, Gradient Descent, and Deep Reinforcement Learning [arifuzzaman2023falcontpds, marlin, jamil2023drl]. These methods offer high utilization with reduced overhead by periodically collecting socket statistics during transfers. The black-box design allows for instant deployment across various networks, as they don’t rely on historical data. These studies finds out that, Concurrency is the most impactful parameter for data transfers. As a result, these solutions narrow the problem to optimizing the number of Concurrency, which can offset suboptimal settings in other parameters. For instance, on a 100 Gbps link, an optimally tuned connection might achieve 10 Gbps, so the solution would require 10 parallel connections. But if TCP buffer size value is suboptimal, throughput per connection could drop to 1 Gbps, prompting the optimizer to increase concurrency to 100. The optimizer can very quickly, often as little in 30 seconds, find this target value due to reduced search space. While these solutions now can easily achieved high network utilization, the increasing concurrency values might significantly degrades the performance of the existing systems and I/O processes as demonstrated in Figure LABEL:fig:cc_impacts. Although extending these solutions to multi-parameter optimization is fairly straightforward, they often perform much worse because the optimizer spends significant time searching for the optimal configuration due to exponentially increasing search space.
III LDM: Systems Design
@@
III-A Handling Transfer Parameters in LDM
@@ III-B Utility Function
III-C Offline Network Simulator
To employ a PPO agent as the optimizer, we must first train it. Training such an agent typically requires thousands of episodes, each consisting of multiple steps. This process becomes especially challenging when the reward for an action cannot be obtained immediately. In our case, at least three seconds of transfer data must be observed to compute the parameters of the utility function for a single step. As a result, one episode may take 30 seconds to one minute. Even under optimistic assumptions, no more than 3,000 episodes could be completed in a day—insufficient for training a PPO agent effectively. Moreover, in many practical scenarios, intended transfers may last less than one hour, making it impractical to dedicate a full day solely for training.
@@ III-C Offline Network Simulator
The simulator, described in Algorithm LABEL:algo:simulator, models all essential aspects of data transfer in a simplified environment. It assumes an infinite pool of files. The files are no larger than the predefined chunk size to ensure parallelism is respected. The simulation is initialized with four parameters: network bandwidth, maximum throughput per thread, initial concurrency, and simulation duration. These values define the environment in which the agent interacts. Whenever the agent predicts a new action, the simulator triggers a call to the GET_UTILITYGET\_UTILITY method. Instead of instantiating actual concurrent transfers, the simulator emulates them using queues. Initially, the queue is populated with the current concurrency level. Each element of the queue, representing one concurrency thread, contains the current time and the remaining bandwidth available for that thread during a one-second interval. Processing proceeds by popping elements from the queue and invoking the TASK method with the time and per-thread bandwidth. This method emulates the time required to send a chunk and returns the next time the thread will begin sending as well as its updated remaining bandwidth. If the next start time exceeds the probing window, the thread is reinserted into the queue. After the queue is fully processed, the throughput is normalized to compute the reward, and the internal state of the simulator is updated accordingly.
III-D PPO Agent Architecture
@@ III-E1 Exploration and Logging
III-E2 Training in the Simulated Environment
@@
III-E3 Adaptive Concurrency in Production
@@ IV Evaluation
In this section, we first discuss how offline training significantly accelerates the agent training process compared to its predecessor, DRL. We then demonstrate the impact of pipelining and parallelism, alongside concurrency, in improving data transfer performance.
IV-A Training in the Simulated Environment
After the logging and exploration phase, we proceed to the offline training of the agent within a simulated environment. We conducted experiments under different scenarios to estimate the time required to satisfy the convergence criteria described in Section III. Across 15 experiments, the mean number of episodes to convergence was 8,525 with a standard deviation of 5,065. In terms of wall-clock time, convergence required an average of 156 seconds with a standard deviation of 88 seconds. The maximum observed time was 222 seconds, while the minimum was 33 seconds.
@@ IV-A Training in the Simulated Environment
TABLE I: Dataset Description
@@ IV-B Experimental Results
We demonstrated the results through experiments conducted on Fabric [fabric-2019]. To emulate production-grade conditions, we utilized machines from two geographically distinct sites: one located in New York, USA (NEWY) and the other at CERN, Switzerland. Both machines were provisioned with 8 CPU cores, 64 GB of RAM, a Dell Express Flash P4510 1 TB SFF SSD, and an NVIDIA Mellanox ConnectX-6 NIC. While the physical port supports a line rate of 100 Gbps, in practice, we observed a maximum transfer speed of 58 Gbps.
@@
TABLE II: End-to-End Transfer Speed Comparison Across Datasets (Gbps)
@@
IV-B1 Effectiveness of Agent-Based Optimization
@@
To highlight the effectiveness of agent-based methods, we limited the per-concurrency network bandwidth to 3 Gbps. This setup allows us to observe how early convergence and stability influence throughput. The results are summarized in Table II, which reports the mean values along with their corresponding standard deviations.
@@
Among adaptive methods, DRL achieves nearly the same throughput as LDM, but it does so using 33% more concurrency, thereby validating our initial hypothesis. Falcon and GD (Heuristic) perform worse because online optimization requires additional time to identify the concurrency value that maximizes throughput. As demonstrated in Figure 3, Falcon and GD (Heuristic) only reached their mean throughput after 46 seconds and 40 seconds, respectively, compared to just 3 seconds and 6 seconds for DRL and LDM. This figure illustrates the best runs of each method in terms of throughput.
@@
IV-B2 Handling Large Files with Parallelism
We designed Dataset B to include files that would take substantial time to transfer if not split into chunks. The per-concurrency bandwidth was throttled to 3 Gbps. Specifically, the dataset consists of four 55 GB files along with 804 files of size 1 GB. If the large files were to start simultaneously at the beginning, transferring a single 55 GB file with one thread would take approximately 55 GB * 8 / 3 Gbps $\approx$ 147 seconds even under maximum utilization of that thread. To highlight the necessity of parallelism, we deliberately structured the transfer queue such that the large files were not initiated at the same time. Methods with pipelining can efficiently handle these cases by chunking the large files and distributing them across multiple threads. This phenomenon is illustrated in Figure 4. The figure shows a steep drop in throughput for Falcon and DRL, both of which lack parallelism. Their throughput curves also exhibit a staircase-like pattern at the tail end. In DRL, one large file was completed before the drop, resulting in steps at 9, 6, and 3 Gbps. In Falcon, the steps occurred at 12 and 6 Gbps, indicating that two large files were completed in close succession.
This bottleneck directly impacts overall throughput, as reported in Table II. In this scenario, LDMclearly outperforms all baselines, achieving approximately 73% higher throughput than the state-of-the-art DRL.
IV-B3 Effect of Pipelining on Small Files
Dataset C is designed to demonstrate the effect of pipelining. This dataset consists of 238,770 files with sizes ranging from 100 KB to 1 MB. In the earlier cases, the lack of pipelining did not significantly degrade performance because the number of socket closures and reconnections was relatively small. However, with approximately fifty times more files than Dataset A, the overhead of repeatedly closing and reopening sockets becomes the dominant bottleneck. This effect is evident in the performance comparison shown in Table II. Interestingly, Globus outperforms Falcon and DRL in this scenario. Its fixed concurrency setting provides an advantage, since dynamically adjusting concurrency introduces additional socket operations. In contrast, methods that integrate pipelining are able to overcome this bottleneck and clearly outperform the others. Nevertheless, the performance of GD and LDM is somewhat slower than in other datasets, primarily due to the repeated encoding and decoding of file headers during transfers.
IV-B4 Mixed File Transfers Without Bandwidth Throttling

### v2

@@ II Related Work & Motivation
Trade-off Between Performance and Overhead: Both heuristic and supervised machine learning models estimate parameter values before data transfers begin, using fixed configurations throughout the entire process. This static approach prevents adaptation to changing network conditions during long-running transfers. Additionally, these solutions lack a back-off mechanism, meaning that large parameter values may overwhelm systems in various production environments. For example, if $x$ parallel TCP connections are required to fully utilize the available bandwidth, assigning $x$ connections to every new transfer becomes inefficient when multiple concurrent transfers are initiated. In a scenario with $N$ concurrent transfers, $Nx$ connections would be created, leading to diminished throughput per connection (by a factor of $N$), but significantly increased overhead. A more efficient alternative would be to allocate fewer connections, say, 10% of $x$ – to each new transfer, under the assumption that overlapping transfers or other processes will compete for I/O resources. For instance, Globus Online, a widely adopted data transfer service, allows tuning of three key parameters: pipelining, parallelism, and concurrency. However, in production environments, system administrators set these values heuristically and often very conservatively to minimize I/O and network contention. While this approach helps balance overhead with resource utilization, it frequently leaves networks underutilized. As a result, static configurations present a dilemma: either prioritize utilization at the cost of high overhead, or reduce overhead and risk underutilizing the network.
@@
Simplified Architecture Overwhelm Systems: The most recent studies also have implemented adaptive approaches using online black-box optimization algorithms such as Bayesian methods, Gradient Descent, and Deep Reinforcement Learning [arifuzzaman2023falcontpds, marlin, drl]. These methods offer high utilization with reduced overhead by periodically collecting socket statistics during transfers. The black-box design allows for instant deployment across various networks, as they don’t rely on historical data. These studies finds out that, Concurrency is the most impactful parameter for data transfers. As a result, these solutions narrow the problem to optimizing the number of Concurrency, which can offset suboptimal settings in other parameters. For instance, on a 100 Gbps link, an optimally tuned connection might achieve 10 Gbps, so the solution would require 10 parallel connections. But if TCP buffer size value is suboptimal, throughput per connection could drop to 1 Gbps, prompting the optimizer to increase concurrency to 100. The optimizer can very quickly, often as little in 30 seconds, find this target value due to reduced search space. While these solutions now can easily achieved high network utilization, the increasing concurrency values might significantly degrade the performance of the existing systems and I/O processes. Although extending these solutions to multi-parameter optimization is fairly straightforward, they often perform much worse because the optimizer spends significant time searching for the optimal configuration due to exponentially increasing search space.
III Understanding The Parameters
The fundamental challenge in maximizing data transfer throughput over High-Performance Networks (HPNs) is an optimization problem constrained by the Bandwidth-Delay Product (BDP) and the limitations of the TCP protocol. To fully utilize the available bandwidth capacity $Bandwidth_{path}$, the aggregate amount of unacknowledged data in flight must equal the BDP of the network path.
$BDP=Bandwidth_{path}\times RTT$ (1)
For a single TCP stream, the throughput $T_{single}$ is bounded by the sender’s congestion window ($W$) and the Round Trip Time ($RTT$): $T_{single}\leq W/RTT$. In modern wide-area networks (WANs), the theoretical BDP often exceeds the operating system’s maximum allowed TCP buffer size ($W_{max}$), meaning $W_{max}\ll BDP$. Consequently, a single stream, or even a naive combination of streams, cannot saturate the link, leaving valuable bandwidth wasted.
To achieve $T_{aggregate}\approx Bandwidth_{path}$, the application layer must manipulate three distinct parameters: Pipelining, Concurrency, and Parallelism, to ensure the summation of all active congestion windows approaches the BDP: $\sum W_{i}\approx BDP$.
Figure 1: Visual comparison of data transfer behaviors under different optimization strategies. Pipelining eliminates control-channel idle gaps, concurrency transfers multiple files simultaneously, and parallelism splits large files. LDM combines all three to maximize throughput.
III-A Pipelining: Necessity for The Small Files
When transferring datasets containing large numbers of small files, the primary bottleneck shifts from bandwidth to latency. For each file transfer, standard protocols undergo a control channel exchange to open the data channel, followed by the TCP 3-way handshake and the slow-start phase. If the transmission time of the file is shorter than the control overhead (typically >1>1 RTT), the data channel remains idle for significant periods, as shown in the "No Optimization" column of Figure 1.
Pipelining addresses this by treating a sequence of small files as a continuous stream, sending transfer commands back-to-back without waiting for intermediate acknowledgments. The impact of this is quantified in Figure 2. Without pipelining (Green and Red lines), the TCP connection is torn down and re-established for every file, causing the congestion window to reset repeatedly. Pipelining prevents the TCP window size from shrinking to zero due to idle data channel time. As seen in Figure 2 (Blue and Orange lines), pipelining maintains a persistent data channel, allowing the congestion window to grow continuously until it reaches the network limit, regardless of individual file sizes.
Figure 2: Growth of the TCP congestion window over time for 1 MB and 32 MB files, with and without pipelining in a 35 Gbps - 67 ms RTT testbed.
III-B Concurrency: Necessary but Insufficient
Concurrency ($cc$) increases the number of simultaneous file transfers, allowing the application to aggregate the throughput of multiple TCP streams: $T_{agg}=\sum_{i=1}^{cc}T_{i}$. This is the primary mechanism for filling the BDP when $W_{max}$ limits individual streams.
Figure 3 demonstrates the linear scaling provided by concurrency. By adding more concurrent flows, the total volume of data in flight increases, effectively filling the available bandwidth pipe. But unnecessarily high concurrency can introduce overheads such as network congestion, packet loss, and high I/O contention, which ultimately reduce overall throughput. This is why the throughput begins to decline after the concurrency level exceeds 8.
However, if a dataset contains a wide range of file sizes, concurrency alone cannot fully optimize the transfer. As shown in the "Concurrency" column of Figure 1, multiple files may be transferred at once, but overheads begin to dominate when the files are much smaller than the BDP.
Figure 3: Impact of increasing concurrency on aggregate throughput. Throughput scales with the number of concurrent transfers until system and network bottlenecks begin to limit additional gains.
III-C Parallelism: Partitioning Large Files
While concurrency allows multiple files to share the link, it is strictly bounded by the file count ($N$). If the dataset contains a few very large files ($N<CC_{optimal}$), concurrency alone leaves the network underutilized because there are not enough distinct files to create the necessary number of streams. Parallelism improves throughput by splitting large files into multiple chunks and transferring these chunks in parallel. Parallelism is advantageous specifically when the system buffer size is smaller compared to the BDP.
Figure 4 demonstrates this effect. When transferring three 2 GB files without parallelism, throughput increases only up to a concurrency level of three. Beyond this point, additional threads have nothing to do, and the throughput plateaus at roughly 4.3 Gbps. By introducing parallelism, each file is divided into multiple chunks (such as P1 and P2 in Figure 1), effectively creating additional "virtual" files. These chunks allow the system to activate more streams, increase the aggregate congestion window, and approach linear throughput scaling. This behavior is visible in the parallelized portion of Figure 4, where throughput continues to rise as concurrency increases because the dataset now contains many parallel chunks rather than only three files.
Figure 4: Throughput measured while increasing concurrency for a dataset containing three 2 GB files without parallelism.
IV LDM: Systems Design
@@
IV-A Handling Transfer Parameters in LDM
@@
chunksizechunksize is a tunable parameter. If it is set too small relative to the maximum achievable congestion window per stream, each chunk re-enters slow start, creating additional overhead. LDM mitigates this effect through infinite pipelining, which keeps connections warm and reduces the cost of repeated ramp-up, although splitting and merging still introduce their own overhead. Conversely, if the chunk size is too large, streams can experience tail latency because some streams finish early while others remain busy with large residual chunks. In practice, we select a chunk size that balances these effects and prevents pathological fragmentation. In most cases, choosing a chunk size of roughly 10 % of the network bandwidth (or link capacity) is a safe option because even when tail latency occurs, the remaining chunks complete quickly and do not meaningfully delay the overall transfer.
@@ IV-B Utility Function
IV-C Offline Network Simulator
To employ a PPO agent as the optimizer, we must first train it. Training such an agent typically requires thousands of episodes, each consisting of multiple steps. This process becomes especially challenging when the reward for an action cannot be obtained immediately. In our case, at least three seconds of transfer data must be observed to compute the parameters of the utility function for a single step. As a result, one episode may take 30 seconds to one minute. Even under optimistic assumptions, no more than 3,000 episodes could be completed in a day, which is insufficient for training a PPO agent effectively. Moreover, in many practical scenarios, intended transfers may last less than one hour, making it impractical to dedicate a full day solely for training.
@@ IV-C Offline Network Simulator
Algorithm 1 I/O and Network Dynamics Simulator
1:Initialization: Throughputs per thread TPTTPT, network bandwidth, initial concurrency, and simulation duration $T_{\text{end}}$.
2:function Task($t$, rem_bw_threadrem\_bw\_thread)
3:  $noise\leftarrow random(0,0.8)*MaxFileChunkSize$
4:  $transferSize\leftarrow MaxFileChunkSize-noise$
5:  $throughput\_increase\leftarrow 0$
6:  Task duration, $d_{task}\leftarrow 0$
7:  Compute throughput increase considering overall bandwidth and remaining bandwidth of thread, rem_bw_threadrem\_bw\_thread.
8:  Compute $d_{task}$ according to TPTTPT.
9:  Update network throughput and rem_bw_threadrem\_bw\_thread
10:  $t_{\text{next}}\leftarrow t+d_{task}+\epsilon$
11:  return $t_{\text{next}}$, rem_bw_threadrem\_bw\_thread
12:end function
13:function Get_Utility(new_concurrency)
14:  Reset throughput counter.
15:  Schedule initial tasks and push to transfer queue for each thread in new_threads with $t=0$ and rem_bw_thread=TPTrem\_bw\_thread=TPT
16:  while the task queue is not empty do
17:   Pop (t,rem_bw_thread)(t,rem\_bw\_thread) from the queue.
[…]
19:   if $t_{\text{next}}<T_{\text{end}}$ then
20:     Add $(t_{\text{next}},rem\_bw\_thread)$ to the queue.
21:   end if
22:  end while
23:  Normalize throughputs by their finish times.
24:  Compute reward
25:  Update the internal simulator state.
26:  return reward and other necessary information.
27:end function
The simulator, described in Algorithm 1, models all essential aspects of data transfer in a simplified environment. It assumes an infinite pool of files. The files are no larger than the predefined chunk size to ensure parallelism is respected. The simulation is initialized with four parameters: network bandwidth, maximum throughput per thread, initial concurrency, and simulation duration. These values define the environment in which the agent interacts. Whenever the agent predicts a new action, the simulator triggers a call to the GET_UTILITYGET\_UTILITY method. Instead of instantiating actual concurrent transfers, the simulator emulates them using queues. Initially, the queue is populated with the current concurrency level. Each element of the queue, representing one concurrency thread, contains the current time and the remaining bandwidth available for that thread during a one-second interval. Processing proceeds by popping elements from the queue and invoking the TASK method with the time and per-thread bandwidth. This method emulates the time required to send a chunk and returns the next time the thread will begin sending as well as its updated remaining bandwidth. If the next start time exceeds the probing window, the thread is reinserted into the queue. After the queue is fully processed, the throughput is normalized to compute the reward, and the internal state of the simulator is updated accordingly.
IV-D PPO Agent Architecture
@@ IV-E1 Exploration and Logging
IV-E2 Training in the Simulated Environment
@@
IV-E3 Adaptive Concurrency in Production
@@ V Evaluation
The goal of this evaluation is to demonstrate the effectiveness, robustness, and practicality of LDM across a wide range of real-world data transfer scenarios. We first quantify the end-to-end performance gains achieved by LDM on diverse datasets and network conditions. We then evaluate the generalizability of the system across different TCP congestion control algorithms, followed by an analysis of stability and fairness when multiple transfers share the network. To understand the contribution of each subsystem, we conduct a detailed study of pipelining, parallelism, and concurrency in isolation and in combination. Finally, we validate key design choices, including the use of a continuous state space for the reinforcement learning controller. Together, these results provide a comprehensive assessment of LDM and establish its suitability for high-performance data transfer environments. In the experiments described in this section, we used $K=1.02$, $B=10$, and chunksize=512MBchunksize=512~MB.
TABLE I: Network Testbeds Used in Experiments
Testbed Site 1 Site 2 Observed Bandwidth RTT
Fabric NEWY CERN 45 Gbps 85 ms
Fabric MICH MASS 35 Gbps 33 ms
CloudLab Clemson Utah 1 Gbps 67 ms
V-A Experimental Setup
Our evaluation is carried out on two widely used research testbeds, summarized in Table I. The primary experiments are conducted on the Fabric testbed [fabric-2019] across two geographically distant sites: NEWY (New York, USA) and CERN (Geneva, Switzerland). Although the underlying Fabric links for these sites are provisioned at 100 Gbps and 50 Gbps respectively, Fabric operates as a shared experiment network where Layer 3 routed paths are shared among multiple experiments. As a result, the maximum stable throughputs observed end-to-end are lower than the physical link capacities, and the values reported in Table I reflect these empirically measured stable rates.
Both endpoints on Fabric are provisioned with 8 CPU cores, 64 GB of RAM, a Dell Express Flash P4510 1 TB NVMe SSD, and an NVIDIA Mellanox ConnectX-6 adapter. On the NEWY–CERN path, although the NIC supports a 100 Gbps port rate, the highest stable throughput achievable in practice is approximately 45 Gbps.
To broaden the evaluation across different bandwidth-delay regimes, we also conduct experiments on a second Fabric path between MICH and MASS, which offers a stable throughput of 35 Gbps with a 33 ms RTT. Additionally, to assess LDM under constrained network conditions, we include experiments on CloudLab between Clemson and Utah, which provides a 1 Gbps link with a 67 ms RTT. Together, these paths span a wide range of bandwidth-delay products, enabling a comprehensive assessment of LDM across both high-performance and limited-bandwidth environments.
V-B Training in the Simulated Environment
After the logging and exploration phase, we proceed to the offline training of the agent within a simulated environment. We conducted experiments under different scenarios to estimate the time required to satisfy the convergence criteria described in Section IV. Across 15 experiments, the mean number of episodes to convergence was 8,525 with a standard deviation of 5,065. In terms of wall-clock time, convergence required an average of 156 seconds with a standard deviation of 88 seconds. The maximum observed time was 222 seconds, while the minimum was 33 seconds.
@@ V-B Training in the Simulated Environment
The simulator preserves the same observable signals as the real network. In practice, it proved highly representative: after deploying the policy and running approximately one hour of online finetuning ($\approx 120$ episodes), the average concurrency changed by only 1 %. This small adjustment shows that the policy learned in simulation already tracks the real optimal concurrency behavior over time. This negligible improvement makes it impractical for real deployments and is therefore excluded from the final solution.
TABLE II: Dataset Description
@@ V-C End-to-End Performance on Diverse Datasets
We evaluated LDMusing the two Fabric testbeds (NEWY–CERN and MICH–MASS) described earlier to capture performance across high-latency and moderate-latency wide-area links. We used BBR as the primary congestion control algorithm because it is specifically designed to operate efficiently in high-bandwidth environments that mirror our target use cases.
@@
TABLE III: End-to-End Transfer Speed Comparison Across Datasets (Gbps) (Fabric Testbed; NEWY-CERN; BBR congestion Control)
@@
TABLE IV: End-to-End Transfer Speed Comparison Across Datasets (Gbps) (Fabric Testbed; MICH-MASS; BBR congestion Control)
Model Dataset A Dataset B Dataset C Dataset D
Globus 6.2 $\pm$ 0.1 32.0 $\pm$ 0.0 3.1 $\pm$ 0.3 32.0 $\pm$ 0.0 2.9 $\pm$ 0.2 64.0 $\pm$ 0.0 11.3 $\pm$ 0.5 64.0 $\pm$ 0.0
Falcon 19.7 $\pm$ 0.9 13.6 $\pm$ 5.3 9.4 $\pm$ 0.9 7.3 $\pm$ 5.2 0.96 $\pm$ 0.02 16.2 $\pm$ 4.8 9.5 $\pm$ 1.0 17.8 $\pm$ 5.9
DRL 38.5 $\pm$ 0.1 30.3 $\pm$ 1.2 12.4 $\pm$ 0.01 30.3 $\pm$ 1.2 0.98 $\pm$ 0.01 15.3 $\pm$ 0.6 10.7 $\pm$ 0.8 20.7 $\pm$ 0.9
LDM-GD 24.6 $\pm$ 0.2 14.9 $\pm$ 5.6 24.5 $\pm$ 0.6 15.8 $\pm$ 6.2 7.7 $\pm$ 0.4 2.2 $\pm$ 0.8 22.2 $\pm$ 0.2 12.2 $\pm$ 3.6
LDM 34.1 $\pm$ 0.3 22.9 $\pm$ 0.8 35.1 $\pm$ 0.4 27.4 $\pm$ 1.0 7.9 $\pm$ 0.4 2.6 $\pm$ 0.9 26.3 $\pm$ 1.2 15.2 $\pm$ 1.8
TABLE V: End-to-End Transfer Speed Comparison Across Datasets (Gbps) (Fabric Testbed; MICH-MASS; CUBIC Congestion Control)
Model Dataset A Dataset B Dataset C Dataset D
Falcon 15.6 $\pm$ 0.4 12.9 $\pm$ 4.2 8.2 $\pm$ 0.2 7.3 $\pm$ 5.4 1.0 $\pm$ 0.0 23.7 $\pm$ 8.1 5.2 $\pm$ 1.5 11.9 $\pm$ 6.3
DRL 21.6 $\pm$ 2.2 29.3 $\pm$ 1.4 9.2 $\pm$ 0.1 13.3 $\pm$ 0.8 1.25 $\pm$ 0.0 28.7 $\pm$ 1.1 7.4 $\pm$ 0.3 17.3 $\pm$ 0.7
LDM-GD 18.3 $\pm$ 0.7 17.1 $\pm$ 6.4 19.9 $\pm$ 1.3 16.43 $\pm$ 6.3 7.9 $\pm$ 1.0 1.7 $\pm$ 0.8 14.9 $\pm$ 0.9 7.8 $\pm$ 3.6
LDM 23.8 $\pm$ 1.5 27.1 $\pm$ 1.7 21.1 $\pm$ 0.7 19.5 $\pm$ 1.9 7.4 $\pm$ 0.2 3.2 $\pm$ 0.1 19.7 $\pm$ 0.6 10.5 $\pm$ 0.6
V-C1 Effectiveness of Agent-Based Optimization
@@
To highlight the effectiveness of agent-based methods, we limited the per-concurrency bandwidth to 3 Gbps on the NEWY–CERN path and 1.5 Gbps on the MICH–MASS path. This setup allows us to observe how early convergence and stability influence throughput.
@@
Among adaptive methods, DRL achieves nearly the same throughput as LDM, but it does so using $\approx$ 33% more concurrency in both the testbeds, thereby validating our initial hypothesis. In the MICH–MASS path, DRL performs better than LDM, and this difference can be explained using the bandwidth–delay product (BDP). On the NEWY–CERN path, the maximum achievable congestion window for a single stream is approximately 3000 Mbps (throttled speed) multiplied by 85 ms RTT, which equals 255 Mb, or about 32 MB. On the MICH–MASS path, the corresponding value is 1500 Mbps multiplied by 33 ms RTT, which equals 49.5 Mb, or about 6 MB. In the first case, the file size is roughly eight times larger than the maximum congestion window, while in the second case it is more than forty times larger. As a result, on the MICH–MASS path, the slow-start overhead becomes negligible compared to the additional processing overhead introduced by infinite pipelining and parallelism, making DRL appear faster. However, such homogeneous file distributions are uncommon in real scientific workloads, and the advantage does not generalize beyond this specific scenario.
Falcon and GD (Heuristic) perform worse because gradient descent-based online optimization requires additional time to identify the concurrency value that maximizes throughput. As demonstrated in Figure 7, Falcon and GD (Heuristic) only reached their mean throughput after 46 seconds and 40 seconds, respectively, compared to just 3 seconds and 6 seconds for DRL and LDM. This figure illustrates the best runs of each method of the NEWY-CERN path in terms of throughput.
@@
V-C2 Handling Large Files with Parallelism
We designed Dataset B to include files that would take substantial time to transfer if not split into chunks. The per-concurrency bandwidth was throttled to 3 Gbps on the NEWY–CERN path and 1.5 Gbps on the MICH–MASS path. Specifically, the dataset consists of four 55 GB files along with 804 files of size 1 GB. If the large files were to start simultaneously at the beginning, transferring a single 55 GB file with one thread would take approximately 55 GB * 8 / 3 Gbps $\approx$ 147 seconds for the NEWY-CERN path and 55 GB * 8 / 1.5 Gbps $\approx$ 294 seconds for the MICH-MASS path, even under maximum utilization of that thread. To highlight the necessity of parallelism, we deliberately structured the transfer queue such that the large files were not initiated at the same time. Methods with pipelining can efficiently handle these cases by chunking the large files and distributing them across multiple threads. This phenomenon is illustrated for the NEWY-CERN path in Figure 8. The figure shows a steep drop in throughput for Falcon and DRL, both of which lack parallelism. Their throughput curves also exhibit a staircase-like pattern at the tail end. In DRL, one large file was completed before the drop, resulting in steps at 9, 6, and 3 Gbps. In Falcon, the steps occurred at 12 and 6 Gbps, indicating that two large files were completed in close succession. It can be seen that DRL achieves higher throughput than LDMuntil the medium-sized files finish transferring. This occurs because the processing overhead of LDM, as explained in the previous section, outweighs its benefits when sufficient medium files, which are much larger than the maximum possible congestion window, are still available.
This bottleneck directly impacts overall throughput, as shown in Table III and Table IV. In this scenario, LDMclearly outperforms all baselines, achieving approximately 73% higher throughput than DRL on the NEWY–CERN path and 273% higher throughput on the MICH–MASS path. The improvement appears larger on MICH–MASS because the bottleneck files operate at roughly half the per-stream speed (as we bottlenecked) compared to NEWY–CERN, which amplifies the benefits of LDM’s optimizations.
V-C3 Effect of Pipelining on Small Files
Dataset C is designed to demonstrate the effect of pipelining. This dataset consists of 238,770 files with sizes ranging from 100 KB to 1 MB. In the earlier cases, the lack of pipelining did not significantly degrade performance because the number of socket closures and reconnections was relatively small. However, with approximately fifty times more files than Dataset A, the overhead of repeatedly closing and reopening sockets becomes the dominant bottleneck. This effect is evident in the performance comparison shown in Table III and Table IV. Interestingly, Globus outperforms Falcon and DRL in this scenario. Its fixed concurrency setting provides an advantage, since dynamically adjusting concurrency introduces additional socket operations. In contrast, methods that integrate pipelining are able to overcome this bottleneck and clearly outperform the others. Nevertheless, the performance of LDM-GD and LDMis somewhat slower than in other datasets, primarily due to the repeated encoding and decoding of file headers during transfers.
V-C4 Mixed File Transfers Without Bandwidth Throttling